OpenStudy (anonymous):

If f(x) = -2x + 3 and g(x) = 4x - 3, which is greater, f(5) or g(-2)?

OpenStudy (solomonzelman):

plug in 5 into f(x) for x and plug in -2 into g(x) for x f(x) = -2x + 3 g(x) = 4x - 3 f(5) = -2(5) + 3 g(-2) = 4(-2) - 3 f(5) = -10 + 3 g(-2) = -8 - 3 f(5) = -7 g(-2) = -11 can you tell me which one is greater ?

OpenStudy (anonymous):

yes i just did that, um -7 right? because -7 is greater on a graphing scale.

OpenStudy (anonymous):

just plugged the value of x

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thank you for plugging it in

OpenStudy (anonymous):

in negative smaller nuber has large value

OpenStudy (anonymous):

yes i was womndering that, thank you :)

OpenStudy (anonymous):

@3tettgerty23 lolll.... @SolomonZelman deserve the medal he solved it for u ...plzzz undo n give the medal to him plzzz

OpenStudy (anonymous):

oh dang! i thought you and him were the same people! sorry soloman! i thought i gave you it

OpenStudy (solomonzelman):

It doesn't matter that much, the medal is not the main thing.

OpenStudy (anonymous):

but yeah, i was tlaking to you solomon! your the one who gave me the work to see how to do it. thanks !!

OpenStudy (anonymous):

yes true :)

OpenStudy (solomonzelman):

Alright, you welcome! Good luck with the rest (if you got any) !

OpenStudy (anonymous):

yeah it is not main thing for u or me but for some its like encouragement.... @solomonZelman

OpenStudy (solomonzelman):

yeah ;) I just keep getting medals and my profile looks suspiciously good.

OpenStudy (anonymous):

haha yeah!

OpenStudy (anonymous):

yeah your al level 99 :)

OpenStudy (solomonzelman):

I got that level on my 1st 2 weeks. and I am like when will I get to 100? (ik, never :( )

OpenStudy (anonymous):

lol youll get there believe me! you answer relaly good and your super smart

OpenStudy (anonymous):

but anyways i gotta go back to my other class :(

OpenStudy (solomonzelman):

Well, if satelline73 gets suspended I might get a 100 :)

OpenStudy (solomonzelman):

Suspended for ever... but he won't

OpenStudy (anonymous):

hahaha